OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form.

OpenStudy (anonymous):

y=m(x-x1)+y1 Plug in your given information into the equation

OpenStudy (anonymous):

y - y1 = m(x - x1) <--point slope form using points (2,5) and slope 3 y - 5 = 3(x - 2) <--your answer

OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form

OpenStudy (anonymous):

Standard form is : Ax + By = C First we have to find the slope(m) m = (y2 - y1) / (x2 - x1) set 1 (2,2).....set 2 (6,3) m = (3 - 2)/(6 - 2) m = 1/4 your slope is 1/4 Now we use y = mx + b using either of the points....I will be using (2,2) and the slope(m) is 1/4 y = mx + b 2 = 1/4(2) + b 2 = 1/2 + b 2 - 1/2 = b 4/2 - 1/2 = b 3/2 = b y = 1/4x + 3/2 (now we need to turn it into standard form) to do this, we multiply by common denominator 4 (4)y = (4)1/4x + (4)3/2 4y = x + 6 -x + 4y = 6 (now multiply it by -1 to make x positive) x - 4y = - 6 <-- this is your equation in standard form